Timber Enterprises purchased 69,300 pounds (cost = $727,650) of direct material to be used in the manufacture of the company's only product. According to the production specifications, each completed unit requires six pounds of direct material at a standard cost of $10 per pound. Direct materials consumed by the end of the period totaled 67,900 pounds in the manufacture of 16,800 finished units.

An examination of Timber’s payroll records revealed that the company worked 50,000 labor hours (cost = $675,000) during the period, and specifications called for each completed unit requiring 2.9 hours of labor at a standard cost of $14.0 per hour.

Assume that the company computes variances at the earliest point in time.

Required:

Compute the following variances. (Indicate the effect of each variance by selecting "Favorable" or "Unfavorable". Select "None" and enter "0" for no effect (i.e., zero variance).)

Direct-material price variance
Direct-material quantity variance
Direct-labor rate variance
Direct-labor efficiency variance
Answer #1

Materials Price Variance = (SP-AP) * AQ

= (10 - 727650/69300) * 69300

= 34,650 Unfavorable

Materials Quantity Variance = (SQ-AQ) * SP

= (16800*6 - 67900)*10

= 329,000 Favourable

Labor Rate Variance = (SR-AR) * AH

= (14 - 675000/50,000)*50000

= 25,000 Favourable

Labor efficiency variance = (SH-AH) * SR

= (16800*2.9 - 50,000)*14

= 17,920 Unfavorable

Note : Comment if you face any issues